OpenStudy (superhelp101):

How many moles of glucose (C6H12O6) are in 1.5 liters of a 4.5 M C6H12O6 solution? 4.5 moles 6.0 moles 6.8 moles 3.0 moles

OpenStudy (superhelp101):

@Abhisar

OpenStudy (abhisar):

We did this type of problem yesterday !

OpenStudy (superhelp101):

i might have an idea how to do this 91.0 = mass NaCl x 100/315 mass NaCl = 286.7 g moles NaCl = 286.7/58.44 g/mol=4.90 moles glucose = 1.5 x 4.5 = 6.8 ???

OpenStudy (abhisar):

4.5 M means 4.5 moles per liters. So no of moles in 1.5 liters is simply 1.5*4.5 = 6.75 moles

OpenStudy (superhelp101):

yeah i was rounding

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @superhelp101 i might have an idea how to do this 91.0 = mass NaCl x 100/315 mass NaCl = 286.7 g moles NaCl = 286.7/58.44 g/mol=4.90 moles glucose = 1.5 x 4.5 = 6.8 ??? \(\color{blue}{\text{End of Quote}}\) What is this ???

OpenStudy (superhelp101):

no! did i do it wrong?

OpenStudy (abhisar):

NaCl ??

OpenStudy (superhelp101):

oh where is my brain!!!!!

OpenStudy (abhisar):

In your head !

OpenStudy (superhelp101):

i wish

OpenStudy (abhisar):

It is !

OpenStudy (abhisar):

You got it ?

OpenStudy (superhelp101):

yep thank you!